OpenStudy (anonymous):

determine the values of x using a x and y chart and the equation 2(x-3)^2 i need x and y coordinates

OpenStudy (turingtest):

I assume you mean y=2(x-3)^2 to find the y-intercepts, set x=0: y=2(-3)^2=18 to find x-intercepts set y=0: 0=2(x-3)^2 x=3 the vertex is (3,0)

OpenStudy (anonymous):

sorry i mean f(x)=2(x-3)^2=2x-2 determine the c values that solve the equation

OpenStudy (turingtest):

??????? you mean solve 2(x-3)^2=2x-2 for x and call that number c? what is c? it's not in the equation.

OpenStudy (anonymous):

:D yeah umm the only thing it says is f(x)=2(x-3)^2=2x-2

OpenStudy (turingtest):

well all I can think to do is solve for x, you want me to do that?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

thak you by the way

OpenStudy (anonymous):

Thank* you u been so helpful btw this is when you plug in numbers for x also

OpenStudy (turingtest):

mulptiply out the exponent:\[2(x-3)^2=2x-2=2(x^2-6x+9)\]divide it all by 2: \[x-1=x^2-6x+9\]collect like terms:\[x^2-7x+10=0\]factor:\[(x-5)(x-2)=0\]\[x=\left\{ 2,5 \right\}\]

OpenStudy (anonymous):

Thank you

OpenStudy (turingtest):

I'm not sure what you mean by "this is when you plug in numbers for x", but if you plug in the values above for x you get (0)x(0)=0 so that verifies our work.

OpenStudy (anonymous):

ohhh thank you